Find the complement and supplement of a 38 degree, 20 minute angle.
Answer by Fombitz(32388)

C+38*20'=90*00'
C=89*60'-38*20'
C=51*40'
.
.
.
S+38*20'=180*00'
S=179*60'-38*20'
S=141*40'